# Nondeterministic Polynomial-time Problem Challenge: An Ever-Scaling Reasoning Benchmark for LLMs

### v1

@@ 6 Conclusions
Reasoning stands as the foundational capability of large language models (LLMs). However, the rapid advancement of LLMs’ reasoning abilities has rendered current benchmarks easily crushable and vulnerable to hacking. Therefore, we propose Nondeterministic Polynomial Problem challenge (NPPC), an ever-scaling benchmark that is uncrushable, unhackable, auto-verifiable, and general, designed to evolve alongside LLM advancements. NPPC comprises three core components: i) npgym: a unified framework for generating customizable problem instances across 25 NPC problems with adjustable complexity levels; ii) npsolver: a flexible evaluation interface supporting both online APIs and offline local deployments; iii) npeval: a comprehensive toolkit for the systematic evaluation of LLMs across different problems, including the solution validity, reasoning errors, token efficiency. Our extensive experiments with state-of-the-art LLMs demonstrate that: i) NPPC successfully reduces all models’ performance to below 10% at extreme difficulties, confirming its uncrushable nature, ii) DeepSeek-R1, Claude-3.7-Sonnet, and o1-mini emerge as the most powerful LLMs, with DeepSeek-R1 outperforming others in 7/12 problems, iii) Models exhibit distinct failure patterns, including cascading assumptions, manual computation errors, and reasoning repetition. To the best of our knowledge, NPPC is the first ever-scaling reasoning benchmark of LLMs, serving as the uncrushable and unhackable testbed for LLMs toward artificial general intelligence (AGI).
@@ Appendix A Frequently Asked Questions (FAQs)
A.1 Why Ever-Scaling and Why the Four Properties are Important?
Why Ever-Scaling?
LLMs are advancing at an unprecedented pace, making existing benchmarks obsolete quickly and posing a significant challenge for maintaining reliable evaluation. An ever-scaling benchmark can evolve alongside LLMs, i.e., adapting dynamically to match the development of LLMs. The ever-scaling benchmark can address two core limitations in traditional benchmarks: i) short lifespan, where traditional benchmarks are easily crushed as LLMs rapidly improve, losing their ability to distinguish between models; ii) limited exploitability, where models can hack the answers in static benchmarks through overfitting or finding shortcuts to answers without genuine reasoning.
Why the Four Properties are Important?
The four properties include:
  • Uncrushable (scaling over the complexity): The benchmark can generate problems with continually increasing difficulty, e.g., larger input sizes, stricter constraints, etc. This property can prevent the benchmark from being solved to prevent obsolescence, and mirror the real-world problems, e.g., logistics and chip design, which grow in complexity as systems scale.
  • Unhackable (scaling over the instances): The benchmark can generate infinite unique instances, even at the same complexity level. This unhackable property makes it impossible for LLMs to memorize the answers or simply overfit to patterns in static training data, and it forces LLMs to reason about the underlying logic to ensure the fairness of evaluation.
  • Auto-verifiable (scaling over the scalable oversight): The benchmark provides an automated and cost-effective evaluation without any human intervention, i.e., the solutions can be verified efficiently in polynomial time even for arbitrarily complex problems. This property is critical for large-scale benchmarking as human evaluation is impractical for massive or highly complex benchmarks, therefore, automated verification is necessary for evaluating at scale.
  • General (scaling over the coverage): This property ensures the benchmark to focus on problems with broad applicability to reflect real-world utility and challenges, enabling the progress made on benchmarks to indicate the progress made on real tasks.
@@ A.2 Why Focusing on NP (Specifically NPC) Problems?
Why not P or NP-hard problems?
The problems in P can be solved in polynomial time. If the LLM is equipped with the tools to run the code, it can generate the code to solve these problems. In this case, the benchmark can be easily crushed by the tool using. While for NP-hard problems, especially for the problems that cannot be verified in polynomial time, when the problems become extremely large, we cannot verify the solution efficiently, which may hurt the scaling over complexity.
Why NPC Problems?
NPC problems are the “hardest” problems in NP class and any other NP problems can be reduced to NPC problems in polynomial time. The absence of known polynomial-time algorithms for NPC problems ensures that current benchmarks measuring performance on these problems cannot be trivially dominated through tool using. Furthermore, the polynomial-time verifiability of solutions enables efficient assessment of solutions generated by LLMs or AI agents even for large problem instances.
@@ A.3 Why Not Considering More Complex Test-time Scaling?
The Majority Voting, Best of $N$, and even tools, e.g., domain-specific solvers, can further improve the performance of models [23, 20]. However, these approaches either necessitate multiple forward passes through the language model or incorporate auxiliary components such as reward models or external tools to augment the reasoning process. Our primary objective is to investigate the reasoning capabilities of LLMs and these complex test-time scaling would be beyond the scope of this paper.
A.4 Why More NP-complete Problems are Needed, i.e., Why Not Focusing on 3SAT Only?
3SAT is a classic NPC problem with theoretical completeness, which provides a theoretically rigorous foundation for benchmarking. As an NPC problem, although all NP problems can be reduced to 3SAT, solely relying on reduction to 3SAT is impractical and reasoning benchmarks demand broader diversity for several key reasons:
  • Reduction overhead: The reduction process may incur significant computational overhead. Additional variables and constraints are often introduced when reducing non-trivial NP problems to a specific NP-complete problem, e.g., reducing Traveling Salesman Problem (TSP) to 3SAT requires mapping the structure of the original problem into a Boolean logic expression through an encoding mechanism, which introduces an exponential number of variables and clauses, significantly increasing the computational complexity.
@@
A.5 Discussion about Future Works
Unstoppable RL vs. Ever-Scaling NP Problems. The rapid progress in LLM reasoning capabilities through reinforcement learning (RL) presents an interesting dynamic when considered alongside ever-scaling NPC problems. As models like DeepSeek-R1 and OpenAI o1/o3-mini demonstrate significant reasoning improvements through RL techniques, NPPC provides a counterbalance by offering problems that can continuously scale in difficulty. This creates an adversarial paradigm to drive the AI development: RL improves model reasoning and NPPC scales to maintain challenging.
Multimodal NP Problems. Extending NPPC to the multimodal domains represents a promising direction. Games like StarCraft II, Minesweeper, Pokemon and Super Mario Bros [3], could form the foundation of a multimodal version of NPPC. This would enable testing reasoning capabilities across multiple modalities while maintaining the ever-scaling properties. A multimodal NPPC could evaluate how well models can reason about visual, spatial, and temporal information jointly.
AI Agent. The benchmark could significantly contribute to AI agent development by encouraging tool use for solving increasingly complex NP problems. As the difficulty of problems increases, LLMs will naturally require external tools to manage computational complexity. This creates a natural pathway toward agent capabilities, where models learn to decompose problems and leverage appropriate tools. The code generation already observed in models attempting to solve difficult NPPC problems can be viewed as a form of tool creation, as these generated algorithms can be saved and reused for future problem-solving. This provides a principled way to measure progress in agent development within a well-defined formal framework.
Appendix B Computational Complexity: P, NP and NP-complete
@@
P.
The class P consists of decision problems that can be solved by a deterministic Turing machine in polynomial time. In practical terms, these are problems for which efficient algorithms exist. The time required to solve these problems grows polynomially with the input size ($n$), such as $O(n)$, $O(n^{2})$, or $O(n^{3})$. Examples include sorting, searching in a sorted array, and determining if a number is prime.
NP.
NP contains all decision problems for which a solution can be verified in polynomial time. Every problem in P is also in NP, but NP may contain problems that are not in P. The key characteristic is that if someone gives you a potential solution, you can quickly check whether it’s correct, even if finding that solution might be difficult. Examples include the Boolean satisfiability problem and the Traveling Salesman decision problem.
NP-complete.
NP-complete problems are the “hardest” problems in NP. A problem is NP-complete if: i) It belongs to NP, ii) Every other problem in NP can be reduced to it in polynomial time. This means that if an efficient (polynomial-time) algorithm were found for any NP-complete problem, it could be used to solve all problems in NP efficiently. The first proven NP-complete problem was the Boolean satisfiability problem (SAT). Other examples include the Traveling Salesman Problem, Graph Coloring, and the Knapsack Problem. The question of whether P=NP (whether every problem with efficiently verifiable solutions also has efficiently computable solutions) remains one of the most important open questions in computer science and mathematics.

### v2

@@ 7 Conclusion
We propose Nondeterministic Polynomial Problem challenge (NPPC), an ever-scaling benchmark that is designed to evolve alongside LLM advancements. NPPC comprises three core components: i) npgym: a unified framework for generating customizable problem instances across 25 NPC problems with adjustable complexity levels; ii) npsolver: a flexible evaluation interface supporting both online APIs and offline local deployments; iii) npeval: a comprehensive toolkit for the systematic evaluation of LLMs across different problems, including the solution validity, reasoning errors, token efficiency. Our extensive experiments with state-of-the-art LLMs demonstrate that: i) NPPC successfully reduces all models’ performance to below 10% at extreme difficulties, confirming its uncrushable nature, ii) DeepSeek-R1, Claude-3.7-Sonnet, and o1-mini emerge as the most powerful LLMs, with DeepSeek-R1 outperforming others in 7/12 problems, iii) Models exhibit distinct failure patterns, including cascading assumptions, manual computation errors, and reasoning repetition. To the best of our knowledge, NPPC is the first ever-scaling reasoning benchmark for reliable and rigorous evaluation of the reasoning limits of LLMs and suggesting the further improvements.
Acknowledgments
This work was partially supported by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (Proposal ID: 23-SIS-SMU-037). The work described in this paper was partially supported by a grant from the Innovation and Technology Commission of the Hong Kong Special Administrative Region, China (Project No. GHP/391/22). This work was also partially supported by the Wallenberg AI, Autonomous Systems and Software Program funded by the Knut and Alice Wallenberg Foundation. The computations were enabled by the supercomputing resource Berzelius provided by the National Supercomputer Centre at Linköping University and the Knut and Alice Wallenberg Foundation, Sweden. Any opinions, findings, and conclusions or recommendations expressed in this publications are those of the author(s) and do not necessarily reflect the views of the funding agencies.
@@ Appendix A Frequently Asked Questions (FAQs)
A.1 Why Ever-Scaling and the Four Desiderata?
Why Ever-Scaling? LLMs are advancing at an unprecedented pace, making existing benchmarks obsolete quickly and posing a significant challenge for maintaining reliable evaluation. An ever-scaling benchmark can evolve alongside LLMs, i.e., adapting dynamically to match the development of LLMs. The ever-scaling benchmark can address two core limitations in traditional benchmarks: i) short lifespan, where traditional benchmarks are easily crushed as LLMs rapidly improve, losing their ability to distinguish between models; ii) limited exploitability, where models can hack the answers in static benchmarks through overfitting or finding shortcuts to answers without genuine reasoning.
Why the Four Desiderate are Important? The four desiderata include:
  • Scaling over complexity. The benchmark can generate problems with continually increasing difficulty, e.g., larger input sizes, stricter constraints, etc. This property can prevent the benchmark from being solved to prevent obsolescence, and mirror the real-world problems, e.g., logistics and chip design, which grow in complexity as systems scale. The scaling over complexity implies if the LLMs solve the generated problem instances of the current difficulty level, the benchmarks can generate more difficult problem instances until the reasoning limits of them.
  • Scaling over the instance. The benchmark can generate infinite unique instances, even at the same complexity level. This property makes it impossible for LLMs to memorize the answers or simply overfit to patterns in static training data, and it forces LLMs to reason about the underlying logic to ensure the fairness of evaluation. To mitigate memorization effects, researchers can randomly sample novel problem instances during evaluation to obtain reliable performance metrics.
  • Scaling over oversight. The benchmark provides an automated and cost-effective evaluation without any human intervention, i.e., the solutions can be verified efficiently even for arbitrarily complex problems. This property is critical for large-scale benchmarking as human evaluation is impractical for massive or highly complex benchmarks, therefore, automated verification is necessary for evaluating at scale.
  • Scaling over coverage. This property enables the benchmark to prioritize problems with broad applicability, thereby reflecting real-world utility and challenges. Consequently, advances demonstrated on the benchmark serve as reliable indicators of progress on practical, real-world tasks.
@@ A.2 Why Focusing on NP (Specifically NPC) Problems?
Why not P or NP-hard Problems? Problems in the P complexity class can be solved in polynomial time. When LLMs are equipped with code execution capabilities, they can generate and execute algorithms to solve these problems directly. Consequently, such benchmarks become susceptible to trivial solutions through computational tools rather than genuine reasoning. Conversely, NP-hard problems, particularly those lacking polynomial-time verification procedures, present scalability challenges: as problem instances grow extremely large, efficient solution verification becomes intractable, potentially compromising the benchmark’s ability to scale over complexity and oversight.
Why NPC Problems? NPC problems are the “hardest” problems in NP class and any other NP problems can be reduced to NPC problems in polynomial time. The absence of known polynomial-time algorithms for NPC problems ensures that current benchmarks measuring performance on these problems cannot be trivially dominated through tool using. Furthermore, the polynomial-time verifiability of solutions enables efficient assessment of solutions generated by LLMs or AI agents even for large-scale problem instances.
Real-World Relevance of Large-Scale NPC Problems. The ability to effectively solve large-scale NPC problems holds profound implications across numerous critical domains in modern society. In logistics and supply chain management, vehicle routing problems for major delivery companies like Amazon or FedEx involve optimizing routes for tens of thousands of vehicles across millions of delivery locations daily, while global supply chain optimization can encompass networks with hundreds of thousands of nodes. Computational biology relies heavily on solving NPC problems where protein folding prediction may explore conformational spaces with $10^{100}$ or more possible states, and genome assembly for complex organisms processes billions of DNA fragments. In telecommunications and computer networks, resource allocation and network design problems routinely involve graphs with millions of nodes and edges, where even moderate-sized instances with thousands of variables become computationally prohibitive. Therefore, developing and evaluating methods capable of tackling NPC problems at truly large scales is not merely an academic exercise but a practical necessity with enormous economic and societal stakes. This makes NPPC particularly valuable as a benchmark: by providing an ever-scaling framework that can generate arbitrarily large NP-complete problem instances, NPPC enables rigorous evaluation of whether AI systems can bridge the critical gap between solving toy problems and addressing the massive-scale combinatorial challenges in real-world applications.
@@ A.3 Why Not Considering More Complex Test-time Scaling?
The Majority Voting, Best of $N$, and even tools, e.g., domain-specific solvers, can further improve the performance of models (Parashar et al., 2025; Lin et al., 2025). However, these approaches either necessitate multiple forward passes through the language model or incorporate auxiliary components such as reward models or external tools to augment the reasoning process. Our primary objective is to investigate the reasoning capabilities of LLMs and these complex test-time scaling would be beyond the scope of this paper. We will tackle this in the future work.
A.4 Why Not Focusing on 3SAT Only?
3SAT is a classic NPC problem with theoretical completeness, which provides a theoretically rigorous foundation for benchmarking. As an NPC problem, although all NP problems can be reduced to 3SAT, solely relying on reduction to 3SAT is impractical and reasoning benchmarks demand broader diversity:
  • Reduction overhead: The reduction process may incur significant computational overhead. Additional variables and constraints are often introduced when reducing non-trivial NP problems to a specific NP-complete problem, e.g., reducing Traveling Salesman Problem (TSP) to 3SAT requires mapping the structure of the original problem into a Boolean logic expression through an encoding mechanism, which introduces an exponential number of variables and clauses, significantly increasing the computational complexity and leading to hidden costs.
@@ A.4 Why Not Focusing on 3SAT Only?
A.5 Can Tool Use Crush NPPC?
Tool use represents a significant advancement in LLM capabilities for tackling NPC problems in the NPPC benchmark, yet fundamental computational barriers remain. When equipped with external tools such as code interpreters, symbolic solvers, or verification systems, LLMs would demonstrate substantially improved performance on NPPC tasks by offloading computationally intensive operations and leveraging specialized algorithms. For instance, tools enable models to execute exhaustive search procedures more reliably, verify candidate solutions programmatically, and utilize domain-specific heuristics that would be difficult to implement through pure text generation. However, despite these enhancements, tool-augmented LLMs still cannot fully solve NPC problems in the general case. The core limitation stems from the inherent computational complexity: while tools can accelerate specific subroutines or handle particular problem instances more efficiently, they do not fundamentally alter the exponential worst-case complexity of NPC problems and we do not have the tools which can solve the NPC problems in polynomial time. Moreover, the effectiveness of tool use depends critically on the model’s ability to decompose problems correctly, formulate appropriate tool calls, and reason about the results. These capabilities remain imperfect even in state-of-the-art systems. Thus, while tool integration marks a meaningful step toward more capable problem-solving systems, NPPC would still serve a meaningful benchmark for tool-augmented LLMs.
A.6 Determining the Difficulty Levels
How to Determine the Difficulty Levels? For NPPC, we address this challenge through a two-stage method: we begin with manual configuration of problem parameters based on established computational complexity theory and domain expertise and then the human-configured difficulty levels are further calibrated through systematic empirical testing with state-of-the-art LLMs. Specifically,
  1. 1.
    Human-defined difficulty parameters: We start with interpretable problem parameters. For SAT, this means specifying (num_variables, num_clauses) such as (3, 5), (4, 5), or (100, 100). While we can confidently say (100, 100) is harder than (5, 5), distinguishing between similar configurations like (5, 4) versus (4, 5) is non-trivial.
  2. 2.
    LLM-based calibration: We use model performance for two specific purposes: i) Fine-grained ordering: Sorting problems of similar complexity for clearer visualization (e.g., Figure 5), particularly when human intuition cannot definitively rank them. ii) Range validation: Ensuring difficulty levels fall within a meaningful range, i.e., avoiding settings where all models achieve 100% (too easy) or 0% (too hard) accuracy, as neither scenario effectively benchmarks capabilities.
We want to note that this is not circular reasoning. We are not using LLM performance to define what makes problems hard; rather, we use it as a practical tool for (1) ordering problems within human-defined difficulty ranges and (2) validating that our chosen parameter ranges enable effective differentiation between models. This two-stage approach ensures that problems’ difficulty levels are both theoretically grounded and practically meaningful for evaluating LLM capabilities.
Are the Generated Instances Truly Difficult for LLMs? Yes, our validation process confirms this through multiple measures: i) we observe consistent performance degradation across difficulty levels, indicating that our instances successfully challenge LLM capabilities, ii) different difficulty levels produce distinct failure modes, suggesting that instances test different aspects of reasoning ability, and iii) the difficulty progression holds across multiple LLM architectures, indicating robustness beyond specific model biases. Although our approach is conceptually simple, it can trully generate difficult instances.
Why not Focusing on Hardest Instances?
Our goal is to evaluate general reasoning capabilities rather than exploit specific failure modes. By providing a graduated difficulty spectrum, we can assess reasoning development by tracking how LLM performance scales with problem complexity, identify capability boundaries to determine where different reasoning strategies break down, and support practical applications by focusing on difficulties relevant to real-world scenarios.
Why not Using Traditional Tools, e.g., Z3 (De Moura and Bjørner, 2008)?
The difficulty experienced by traditional symbolic solvers does not necessarily translate to difficulty for LLMs due to fundamental differences in problem-solving approaches. First, traditional solvers use systematic search and logical inference, while LLMs rely on pattern recognition and learned heuristics. Second, problems that are hard for symbolic methods due to search space explosion may be tractable for LLMs through pattern matching, and vice versa, therefore, the relationship between problem size and difficulty differs dramatically between symbolic and neural approaches. Third, traditional tools fail due to computational resource constraints, while LLMs fail due to reasoning limitations or training data gaps. Therefore, LLM-specific calibration is essential to create benchmarks that meaningfully assess the unique capabilities and limitations of LLMs. In examining 25 NPC problems across multiple domains, we observe that problem-specific tools, while potentially effective within their narrow scope, lack the generalizability required for comprehensive evaluation. Therefore, we do not rely on traditional computational tools as the primary metric for establishing problem difficulty levels in LLM evaluation frameworks.
A.7 Selection of Models
Due to the limited budget, we can only select the representative models for the evaluation. Specifically, we choose the two representative offline medium-sized reasoning models, i.e., QwQ-32B and DeepSeek-R1-32B, and online advanced non-reasoning models, i.e., GPT-4o-mini, GPT-4o, Claude-3.7-Sonnet, DeepSeek-V3, DeepSeek-V3-2503, and online reasoning models, i.e., DeepSeek-R1, o1-mini, and o3-mini. For the more recent models, e.g., o3, o4-mini, Gemini 2.5 Pro, Qwen 3, Llama 4, Claude-4, and GPT-5, we will add them in the next update of our benchmark.
A.8 Code and Leaderboard
We provide the screenshot of the leaderboard in Figure 11. The code and leaderboard can be accessed at https://github.com/SMU-DIGA/nppc.
Figure 11: Screenshot of NPPC leaderboard
A.9 Correlation of Performance between NPPC and Other Benchmarks
Model NPPC MMLU GPQA Diamond LMArena
DeepSeek-R1 0.62 90.8 71.5 1396
o3-mini 0.51 84.9 70.6 1348
Claude-3.7-Sonnet 0.49 86.1 68.0 1371
DeepSeek-V3-2503 0.43 88.5 68.4 1392
o1-mini 0.34 85.2 60.0 1336
DeepSeek-V3 0.33 88.5 59.1 1358
QwQ-32B 0.3 - 59.5 1334
GPT-4o 0.24 85.7 49.0 1335
DeepSeek-R1-32B 0.17 87.4 62.1 -
GPT-4o-mini 0.14 82.0 - 1317
Table 5: Model performance over benchmarks.
Figure 12: Correlation between NPPC and established benchmarks.
Figure 13: Correlation matrix of benchmarks.
To validate the effectiveness of NPPC as a reasoning benchmark, we analyze its correlation with three widely-used evaluation metrics: MMLU (Hendrycks et al., 2021), GPQA Diamond (Rein et al., 2024), and LMArena (Chiang et al., 2024). We compute both Pearson correlation coefficient (for linear relationships) and Spearman rank correlation coefficient (for monotonic relationships) across the 10 considered representative models. As shown in Figure 12, NPPC demonstrates strong positive correlations with both GPQA Diamond (Pearson’s $r=0.798$, $p=0.010$; Spearman’s $\rho=0.800$, $p=0.010$) and LMArena ($r=0.821$, $p=0.007$; $\rho=0.833$, $p=0.005$), both statistically significant at the $p<0.05$ level. The correlation with MMLU shows a positive trend but does not reach statistical significance ($r=0.553$, $p=0.123$), likely reflecting MMLU’s broader emphasis on factual knowledge alongside reasoning capabilities. These results provide several important validations. First, the strong correlation with GPQA Diamond, a graduate-level science reasoning benchmark, confirms that NPPC effectively captures complex reasoning abilities required for advanced problem-solving. Second, the strong correlation with LMArena, a human preference-based evaluation platform, demonstrates that NPPC performance aligns well with practical model utility in real-world applications. The correlation matrix in Figure 13 provides a comprehensive view of these relationships.
A.10 Broader Impact Statement
NPPC provides a scalable framework grounded in complexity theory for tracking progress in computational reasoning, which can inform development of AI systems for scientific computing, optimization, and algorithm design. The NPPC benchmark may also raise several important considerations:
  • Computational and Environmental Costs. The ever-scaling nature of NPPC requires substantial computational resources, particularly as problem difficulty increases. Large-scale evaluations may result in significant energy consumption and carbon emissions. Future work should explore methods to maintain rigor while reducing computational costs through efficient subset selection or adaptive evaluation protocols.
  • Risk of Overfitting to Synthetic Tasks. Models could be specifically optimized for NPPC without corresponding improvements in real-world reasoning. This could lead to overfitting to benchmark-specific patterns or exploitation of artifacts rather than genuine algorithmic understanding. NPPC should be viewed as one component of comprehensive evaluation, not a singular optimization target. We discourage training practices that narrowly target NPPC performance at the expense of general capability development, real-world robustness, or safety considerations.
  • Scope and Interpretation of Results. High NPPC performance demonstrates proficiency in structured computational reasoning but should not be interpreted as evidence of general intelligence or real-world reliability. We call on researchers to interpret results within proper context and avoid overstating capabilities based on benchmark performance alone.
Appendix B Computational Complexity: P, NP, NP-complete and NP-hard
@@
P. The class P consists of decision problems that can be solved by a deterministic Turing machine in polynomial time. In practical terms, these are problems for which efficient algorithms exist. The time required to solve these problems grows polynomially with the input size ($n$), such as $O(n)$, $O(n^{2})$, or $O(n^{3})$. Examples include sorting, searching in a sorted array, and determining if a number is prime.
NP. NP contains all decision problems for which a solution can be verified in polynomial time. Every problem in P is also in NP, but NP may contain problems that are not in P. The key characteristic is that if someone gives you a potential solution, you can quickly check whether it’s correct, even if finding that solution might be difficult. Examples include the Boolean satisfiability problem and the Traveling Salesman decision problem.
NP-complete (NPC). NP-complete problems are the “hardest” problems in NP. A problem is NP-complete if: i) It belongs to NP, ii) Every other problem in NP can be reduced to it in polynomial time. This means that if an efficient (polynomial-time) algorithm were found for any NP-complete problem, it could be used to solve all problems in NP efficiently. The first proven NP-complete problem was the Boolean satisfiability problem (SAT). Other examples include the Traveling Salesman Problem, Graph Coloring, and the Knapsack Problem. The question of whether P=NP (whether every problem with efficiently verifiable solutions also has efficiently computable solutions) remains one of the most important open questions in computer science.
NP-hard Problems. A problem is NP-hard if every problem in NP can be reduced to it in polynomial time, but unlike NPC problems, NP-hard problems need not be in NP themselves (i.e., they need not be decision problems or have efficiently verifiable solutions). This makes NP-hard a broader class that includes optimization versions of NPC problems (e.g., finding the minimum vertex cover rather than deciding if one of size $k$ exists), as well as problems strictly harder than NP (e.g., certain problems in PSPACE or EXPTIME). In practice, many real-world applications involve NP-hard optimization problems where the goal is to find optimal or near-optimal solutions rather than simply verify their existence.